# Tight frames over the quaternions and equiangular lines

### v2

@@ 2.3 Bounds on equiangular lines
$P_{j}=v_{j}v_{j}^{*}:v\mapsto v_{j}\langle v,v_{j}\rangle,\qquad j=1,\ldots,n,$
@@
(Five equiangular lines in $\mathbb{H}^{2}$). Fix $10<t<$ 1, and consider the four unit vectors
@@
${1\over\sqrt{2}}\begin{pmatrix}1\cr 1\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr i\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr j\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr k\end{pmatrix},\quad\begin{pmatrix}1\cr 0\end{pmatrix}\ \bigl{(}\hbox{or }\begin{pmatrix}0\cr 1\end{pmatrix}\bigr{)}.$
@@
Another method to obtain tight equiangular lines is via the complementary tight frame. The construction is as follows. Let $G$ be the Gramian of $n>d$ equiangular unit vectors in $\mathbb{F}^{d}$ at an angle $\lambda={n-d\over d(n-1)}\neq 0$, so that $P={d\over n}G$ is an orthogonal projection matrix (Proposition 2.1). The complementary orthogonal projection $Q=I-P$ gives an equiangular tight frame of $n$ vectors for $\mathbb{F}^{n-d}$ with Gramian $G_{c}$ given by
@@
This was obtained by the following general method. The condition $VV^{*}=AI$ for $V$ to be a tight frame is that the rows of $V$ are orthogonal and of equal length, i.e., $V^{*}$ has orthogonal columns of equal length. By using Gram-Schmidt, add orthogonal columns of equal length to obtain $[V^{*},W^{*}]$ a scalar multiple of a unitary matrix. Then $W$ is a tight frame, which is the complement of $V$, since
@@
Above we used the fact that the the rows of the matrix $V$ giving a tight frame are orthogonal. For frames over $\mathbb{C}$ this is equivalent to the rows being orthogonal. For frames over the quaternions, it is necessary to make this distinction. Indeed, there exist unitary matrices (orthogonal columns) whose rows are not orthogonal, e.g.,
@@
By Proposition 2.3, there cannot be six tight equiangular lines in $\mathbb{R}^{4}$ or $\mathbb{C}^{4}$, but they do exist in $\mathbb{H}^{4}$, by taking the complementary tight frame to the six tight equiangular lines in $\mathbb{H}^{2}$ of [ET20].
@@
[…]
@@
${d\lambda\over dn}={d-1\over d(n-1)^{2}}>0,$
@@ 2.4 Equi-isoclinic and equichordal subspaces
For $\mathbb{F}^{d}=\mathbb{R}^{d},\mathbb{C}^{d}$, we have $\mathop{\rm trace}\nolimits(P_{j}P_{k})=\mathop{\rm trace}\nolimits(P_{k}P_{j})\in\mathbb{R}$, and a collection of $r$-subspaces is a said to be equichordal (see [FJMW17]) if the corresponding orthogonal projections satisfy
$\mathop{\rm trace}\nolimits(P_{j}P_{k})=\lambda r,\qquad j\neq k,$
@@
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\lambda r,\quad j\neq k\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\|P_{j}-P_{k}\|_{F}^{2}=2(1-\lambda)r,\quad j\neq k.$ (2.17)
Two $r$-subspaces ${\cal V}_{j}$ and ${\cal V}_{k}$, $j\neq k$, are isoclinic with parameter $0\leq\lambda\leq 1$ (see [LS73], [Hog77]) if the orthogonal projection $P_{jk}$ onto ${\cal V}_{j}+{\cal V}_{k}$ satisfies
@@
$(1-\lambda)P_{j}=(P_{j}-P_{k})^{2}P_{j}\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002ptP_{j}P_{k}P_{j}=\lambda P_{j}.$
@@
[…]
@@
$P_{j}P_{k}P_{j}=\lambda P_{j}\hskip 10.00002pt\Longrightarrow\hskip 10.00002pt\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}P_{j}))=\mathop{\rm trace}\nolimits(\lambda P_{j})=\lambda r.$
@@ 3 From $\mathbb{R}$ to $\mathbb{C}$ and $\mathbb{C}$ to $\mathbb{H}$, and back
$[\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}:v\mapsto\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\qquad\mathop{\rm Re}\nolimits v={v+\overline{v}\over 2},\quad\mathop{\rm Im}\nolimits v={v-\overline{v}\over 2i}.$ (3.19)
Based on a thorough analysis of this, we will then define an analogous map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$. The first subtle point, is that $[\cdot]_{\mathbb{R}}$ maps $k$-dimensional complex-subspaces of $\mathbb{C}^{d}$ to real $(2k)$-dimensional subspaces of $\mathbb{R}^{2d}$. To see why this is, we first calculate the image of a complex scalar multiple $\alpha+i\beta$ of a vector $v=x+iy$
@@
$[(\alpha+i\beta)v]_{\mathbb{R}}=\alpha\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix}+\beta\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}=\alpha[v]_{\mathbb{R}}+\beta[iv]_{\mathbb{R}}.$ (3.20)
@@
$[\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}\}\qquad\hbox{(orthogonal vectors in $\mathbb{R}^{2d}$)}.$
@@
$\sum_{\ell}(\alpha_{\ell}+i\beta_{\ell})v_{\ell}=0\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\sum_{\ell}\Bigl{\{}\alpha_{\ell}\begin{pmatrix}\mathop{\rm Re}\nolimits v_{\ell}\cr\mathop{\rm Im}\nolimits v_{\ell}\end{pmatrix}+\beta_{\ell}\begin{pmatrix}-\mathop{\rm Im}\nolimits v_{\ell}\cr\mathop{\rm Re}\nolimits v_{\ell}\end{pmatrix}\Bigr{\}}=0.$
@@
$=\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Re}\nolimits w\rangle+\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Im}\nolimits w\rangle+i(\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Re}\nolimits w\rangle-\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Im}\nolimits w\rangle),$
@@
$\mathop{\rm Re}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[v]_{\mathbb{R}},[w]_{\mathbb{R}}\rangle_{\mathbb{R}},\quad\mathop{\rm Im}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[v]_{\mathbb{R}},[iw]_{\mathbb{R}}\rangle_{\mathbb{R}},\qquad$ (3.21)
@@
Let $A:\mathbb{C}^{n}\to\mathbb{C}^{m}$ a $\mathbb{C}$-linear map be represented as an $\mathbb{R}$-linear map $[A]_{\mathbb{R}}:\mathbb{R}^{2n}\to\mathbb{R}^{2m}$ under this identification, i.e., $[A]_{\mathbb{R}}:=[\cdot]_{\mathbb{R}}A[\cdot]_{\mathbb{R}}^{-1}$. Then
@@
The usual rules for matrix multiplication follow, e.g, $[A]_{\mathbb{R}}[B]_{\mathbb{R}}=[AB]_{\mathbb{R}}$. One must be careful if a vector $v\in\mathbb{C}^{d}$ is being thought of as a $d\times 1$ matrix, i.e., the linear map $[v]:\mathbb{C}\to\mathbb{C}^{d}:\alpha\mapsto\alpha v$, since $[v]_{\mathbb{R}}\in\mathbb{R}^{2d\times 1}$, $[[v]]_{\mathbb{R}}\in\mathbb{R}^{2d\times 2}$. In particular, the familiar formula $P=vv^{*}$ for the orthogonal projection onto a unit vector $v\in\mathbb{C}^{d}$, is $P=[v][v]^{*}$, which maps as follows
$[P]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]^{*}]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]]_{\mathbb{R}}^{T},\qquad[[v]]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v&-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Im}\nolimits v&\mathop{\rm Re}\nolimits v\end{pmatrix}.$
@@
$[\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{[v]_{\mathbb{R}},[iv]_{\mathbb{R}}\},\qquad[v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\quad[iv]_{\mathbb{R}}=\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}.$
The identification $[\cdot]_{\mathbb{R}}$ preserves various properties of linear maps, see Theorem 3.3. In particular, orthogonal projections map to orthogonal projections, and hence:
@@
We now consider the situation for tight frames, which is somewhat more involved, e.g., a basis for $\mathbb{C}^{d}$ does not correspond to a basis for $\mathbb{R}^{2d}$ (which has twice the dimension). Let $V=V_{1}+iV_{2}$ be the synthesis map for a sequence of vectors $v_{1},\ldots,v_{n}\in\mathbb{C}^{d}$, and $V_{\mathbb{R}}$ be the corresponding map for the sequence $[v_{1}]_{\mathbb{R}},\ldots,[v_{n}]_{\mathbb{R}}\in\mathbb{R}^{2d}$, i.e.,
@@
Let $[\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ be the correspondence (3.19) between $\mathbb{C}^{d}$ and $\mathbb{R}^{2d}$. Then
@@
$\sum_{j}\sum_{k}(z_{j}\overline{z_{k}})^{2}=\Bigl{(}\sum_{j}z_{j}^{2}\Bigr{)}\Bigl{(}\sum_{k}\overline{z_{k}}^{2}\Bigr{)}=\Bigl{|}\sum_{j}z_{j}^{2}\Bigr{|}^{2}=0\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\sum_{j}z_{j}^{2}=0.$
@@
We now give a map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$ that has similar properties to $[\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$. This is based on the following analogue of the polar decomposition for $\mathbb{C}$, the Cayley-Dickson construction, that every quaternion $q\in\mathbb{H}$ can be written uniquely
@@
$[\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}:z+wj\mapsto\begin{pmatrix}z\cr\overline{w}\end{pmatrix},$ (3.29)
@@
$[(z+wj)\alpha]_{\mathbb{C}}=\begin{pmatrix}z\alpha\cr\overline{w}\alpha\end{pmatrix}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix}\alpha=[z+wj]_{\mathbb{C}}\alpha\qquad\forall\alpha\in\mathbb{C}.$
@@
$=(\langle v_{1},w_{1}\rangle+\langle\overline{v_{2}},\overline{w_{2}}\rangle)-(\langle-v_{2},w_{1}\rangle+\langle\overline{v_{1}},\overline{w_{2}}\rangle)j,$
$=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}-\langle[vj]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}j$
@@
$\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}})=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}},\qquad\mathop{\rm Co}\nolimits_{2}(\langle v,w\rangle_{\mathbb{H}})=-\langle[vj]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}.$ (3.30)
@@
$\mathop{\rm Re}\nolimits(\langle v,w\rangle_{\mathbb{H}})=\mathop{\rm Re}\nolimits(\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}}))=\mathop{\rm Re}\nolimits(\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}})=\langle[[v]_{\mathbb{C}}]_{\mathbb{R}},[[w]_{\mathbb{C}}]_{\mathbb{R}}\rangle_{\mathbb{R}}.$ (3.32)
@@
$[v(\alpha+\beta j)]_{\mathbb{C}}=[v\alpha+v\beta j]_{\mathbb{C}}=[v\alpha+vj\overline{\beta}]_{\mathbb{C}}=[v]_{\mathbb{C}}\alpha+[vj]_{\mathbb{C}}\overline{\beta},\qquad\alpha,\beta\in\mathbb{C},$ (3.33)

### v3

@@ 2.3 Bounds on equiangular lines
$P_{j}=v_{j}v_{j}^{*}:v\mapsto v_{j}\langle v_{j},v\rangle,\qquad j=1,\ldots,n,$
@@
(Five equiangular lines in $\mathbb{H}^{2}$). Fix $0<t<1$, and consider the four unit vectors
@@
${1\over\sqrt{2}}\begin{pmatrix}1\cr 1\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr i\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr j\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr k\end{pmatrix},\quad\begin{pmatrix}1\cr 0\end{pmatrix}\ \bigl{(}\hbox{or }\begin{pmatrix}0\cr 1\end{pmatrix}\bigr{)}$. (2.14)
@@
Taking the five lines of (2.14) and their orthogonal complement gives five MUBs (mutually unbiased bases) in $\mathbb{H}^{2}$, which is a tight frame of ten vectors (see [BW25]).
Another method to obtain tight equiangular lines is via the complementary tight frame. The construction is as follows. Let $G$ be the Gramian of $n>d$ equiangular unit vectors in $\mathbb{F}^{d}$ at an angle $\lambda={n-d\over d(n-1)}\neq 0$, so that $P={d\over n}G$ is an orthogonal projection matrix (Proposition 2.1). The complementary orthogonal projection Q=IPQ=I-P gives an equiangular tight frame of $n$ vectors for $\mathbb{F}^{n-d}$ with Gramian $G_{c}$ given by
@@
This was obtained by the following general method. The condition $VV^{*}=AI$ for $V$ to be a tight frame is that the entrywise conjugates of the rows of $V$ are orthogonal and of equal length, i.e., $V^{*}$ has orthogonal columns of equal length. By using Gram-Schmidt, add orthogonal columns of equal length to obtain $V^{*},W^{*}]$ a scalar multiple of a unitary matrix. Then $W$ is a tight frame, which is the complement of $V$, since
@@
Above we used the fact that the columns of the square matrix $\begin{pmatrix}V^{*}&W^{*}\end{pmatrix}$ over $\mathbb{H}$ are orthogonal. For frames over $\mathbb{C}$, this is equivalent to the rows being orthogonal. For frames over the quaternions, it is necessary to make this distinction. Indeed, there exist unitary matrices (orthogonal columns) whose rows are not orthogonal, e.g.,
@@
By Proposition 2.3, there cannot be six tight equiangular lines in $\mathbb{R}^{4}$ or $\mathbb{C}^{4}$, but they do exist in $\mathbb{H}^{4}$, by taking the complementary tight frame to the six tight equiangular lines in $\mathbb{H}^{2}$ of [KF08], [ET20] (obtained independently).
@@
$\lambda=0\quad\hbox{(orthonormal basis)},\qquad\lambda={1\over d^{2}}\quad\hbox{(vertices of a simplex)},$
@@
$\Bigl{(}{\partial\lambda\over\partial n}\Bigr{)}_{d}={d-1\over d(n-1)^{2}}>0,$
@@ 2.4 Equi-isoclinic and equichordal subspaces
For $\mathbb{F}^{d}=\mathbb{R}^{d},\mathbb{C}^{d}$, we have $\mathop{\rm trace}\nolimits(P_{j}P_{k})=\mathop{\rm trace}\nolimits(P_{k}P_{j})\in\mathbb{R}$, and a collection of $r$-subspaces is said to be equichordal (see [FJMW17]) if the corresponding orthogonal projections satisfy
$\langle P_{j},P_{k}\rangle_{F}=\mathop{\rm trace}\nolimits(P_{j}P_{k})=\lambda r,\qquad j\neq k,$
@@
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\lambda r,\quad j\neq k\quad\Longleftrightarrow\quad\|P_{j}-P_{k}\|_{F}^{2}=2(1-\lambda)r,\quad j\neq k.$ (2.18)
Two $r$-subspaces ${\cal V}_{j}$ and ${\cal V}_{k}$, $j\neq k$, are isoclinic with parameter $0\leq\lambda\leq 1$ (see [LS73], [Hog77], [FIJM24]) if the orthogonal projection $P_{jk}$ onto ${\cal V}_{j}+{\cal V}_{k}$ satisfies
@@
$(1-\lambda)P_{j}=(P_{j}-P_{k})^{2}P_{j}\quad\Longleftrightarrow\quad P_{j}P_{k}P_{j}=\lambda P_{j}.$
@@
$P_{j}P_{k}P_{j}=\lambda P_{j}\quad\Longleftrightarrow\quad(V_{j}^{*}V_{k})(V_{k}^{*}V_{j})=\lambda I\quad\Longleftrightarrow\quad V_{k}^{*}V_{j}V_{j}^{*}V_{k}=\lambda I\quad\Longleftrightarrow\quad P_{k}P_{j}P_{k}=\lambda P_{k}.$
@@
$P_{j}P_{k}P_{j}=\lambda P_{j}\quad\Longrightarrow\quad\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}P_{j}))=\mathop{\rm trace}\nolimits(\lambda P_{j})=\lambda r.$
@@ 3 From $\mathbb{R}$ to $\mathbb{C}$ and $\mathbb{C}$ to $\mathbb{H}$, and back
$\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}:v\mapsto\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\qquad\mathop{\rm Re}\nolimits v={v+\overline{v}\over 2},\quad\mathop{\rm Im}\nolimits v={v-\overline{v}\over 2i}.$ (3.20)
Based on a thorough analysis of this, we will then define an analogous map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$. The first subtle point, is that $\cdot$$_{\mathbb{R}}$ maps $k$-dimensional complex-subspaces of $\mathbb{C}^{d}$ to real $(2k)$-dimensional subspaces of $\mathbb{R}^{2d}$. To see why this is, we first calculate the image of a complex scalar multiple $\alpha+i\beta$ of a vector $v=x+iy$
@@
$(\alpha+i\beta)v]_{\mathbb{R}}=\alpha\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix}+\beta\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}=\alpha[v]_{\mathbb{R}}+\beta[iv]_{\mathbb{R}}.$ (3.21)
@@
$\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}\}\qquad\hbox{(orthogonal vectors in $\mathbb{R}^{2d}$)}.$
@@
$\sum_{\ell}(\alpha_{\ell}+i\beta_{\ell})v_{\ell}=0\quad\Longleftrightarrow\quad\sum_{\ell}\Bigl{\{}\alpha_{\ell}\begin{pmatrix}\mathop{\rm Re}\nolimits v_{\ell}\cr\mathop{\rm Im}\nolimits v_{\ell}\end{pmatrix}+\beta_{\ell}\begin{pmatrix}-\mathop{\rm Im}\nolimits v_{\ell}\cr\mathop{\rm Re}\nolimits v_{\ell}\end{pmatrix}\Bigr{\}}=0.$
@@
$=\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Re}\nolimits w\rangle+\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Im}\nolimits w\rangle+i(\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Im}\nolimits w\rangle-\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Re}\nolimits w\rangle),$
@@
$\mathop{\rm Re}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[v]_{\mathbb{R}},[w]_{\mathbb{R}}\rangle_{\mathbb{R}},\quad\mathop{\rm Im}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[iv]_{\mathbb{R}},[w]_{\mathbb{R}}\rangle_{\mathbb{R}},\qquad$ (3.22)
@@
Let $A:\mathbb{C}^{n}\to\mathbb{C}^{m}$ a $\mathbb{C}$-linear map be represented as an $\mathbb{R}$-linear map $A]_{\mathbb{R}}:\mathbb{R}^{2n}\to\mathbb{R}^{2m}$ under this identification, i.e., $A]_{\mathbb{R}}:=[\cdot]_{\mathbb{R}}A[\cdot]_{\mathbb{R}}^{-1}$. Then
@@
The usual rules for matrix multiplication follow, e.g, $A]_{\mathbb{R}}[B]_{\mathbb{R}}=[AB]_{\mathbb{R}}$. One must be careful if a vector $v\in\mathbb{C}^{d}$ is being thought of as a $d\times 1$ matrix, i.e., the linear map $v]:\mathbb{C}\to\mathbb{C}^{d}:\alpha\mapsto\alpha v$, since $v]_{\mathbb{R}}\in\mathbb{R}^{2d\times 1}$, $v]]_{\mathbb{R}}\in\mathbb{R}^{2d\times 2}$. In particular, the familiar formula $P=vv^{*}$ for the orthogonal projection onto a unit vector $v\in\mathbb{C}^{d}$, is $P=[v][v]^{*}$, which maps as follows
$P]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]^{*}]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]]_{\mathbb{R}}^{T},\qquad[[v]]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v&-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Im}\nolimits v&\mathop{\rm Re}\nolimits v\end{pmatrix}$.
@@
$\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{[v]_{\mathbb{R}},[iv]_{\mathbb{R}}\},\qquad[v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\quad[iv]_{\mathbb{R}}=\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}$.
The identification $\cdot$$_{\mathbb{R}}$ preserves various properties of linear maps, see Theorem 3.3. In particular, orthogonal projections map to orthogonal projections, and hence:
@@
We now consider the situation for tight frames, which is somewhat more involved, e.g., a basis for $\mathbb{C}^{d}$ does not correspond to a basis for $\mathbb{R}^{2d}$ (which has twice the dimension). Let $V=V_{1}+iV_{2}$ be the synthesis map for a sequence of vectors $v_{1},\ldots,v_{n}\in\mathbb{C}^{d}$, and $V_{\mathbb{R}}$ be the corresponding map for the sequence $v_{1}]_{\mathbb{R}},\ldots,[v_{n}]_{\mathbb{R}}\in\mathbb{R}^{2d}$, i.e.,
@@
Let $\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ be the correspondence (3.20) between $\mathbb{C}^{d}$ and $\mathbb{R}^{2d}$. Then
@@
$\sum_{j}\sum_{k}(z_{j}\overline{z_{k}})^{2}=\Bigl{(}\sum_{j}z_{j}^{2}\Bigr{)}\Bigl{(}\sum_{k}\overline{z_{k}}^{2}\Bigr{)}=\Bigl{|}\sum_{j}z_{j}^{2}\Bigr{|}^{2}=0\quad\Longleftrightarrow\quad\sum_{j}z_{j}^{2}=0.$
@@
We now give a map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$ that has similar properties to $\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$. This is based on the following analogue of the polar decomposition for $\mathbb{C}$, the Cayley-Dickson construction, that every quaternion $q\in\mathbb{H}$ can be written uniquely
@@
$\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}:z+wj\mapsto\begin{pmatrix}z\cr\overline{w}\end{pmatrix}$, (3.30)
@@
$(z+wj)\alpha]_{\mathbb{C}}=\begin{pmatrix}z\alpha\cr\overline{w}\alpha\end{pmatrix}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix}\alpha=[z+wj]_{\mathbb{C}}\alpha\qquad\forall\alpha\in\mathbb{C}.$
@@
$=\langle v_{1},w_{1}\rangle-j\langle v_{2},w_{2}\rangle j-j\langle v_{2},w_{1}\rangle+\langle v_{1},w_{2}\rangle j,$
$=\langle v_{1},w_{1}\rangle+\langle\overline{v_{2}},\overline{w_{2}}\rangle-(\langle\overline{v_{2}},\overline{w_{1}}\rangle+\langle v_{1},-w_{2}\rangle)j,$
$=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}-\langle[v]_{\mathbb{C}},[wj]_{\mathbb{C}}\rangle_{\mathbb{C}}j$
@@
$\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}})=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}},\qquad\mathop{\rm Co}\nolimits_{2}(\langle v,w\rangle_{\mathbb{H}})=-\langle[v]_{\mathbb{C}},[wj]_{\mathbb{C}}\rangle_{\mathbb{C}}.$ (3.31)
@@
$v(\alpha+\beta j)]_{\mathbb{C}}=[v\alpha+v\beta j]_{\mathbb{C}}=[v\alpha+vj\overline{\beta}]_{\mathbb{C}}=[v]_{\mathbb{C}}\alpha+[vj]_{\mathbb{C}}\overline{\beta},\qquad\alpha,\beta\in\mathbb{C},$ (3.33)